### v1

# Data-driven modeling of multiscale phenomena with applications to fluid turbulence

Abstract
This letter introduces a novel data-driven framework for constructing accurate and general equivariant models of multiscale phenomena which does not rely on specific assumptions about the underlying physics. This framework is illustrated using incompressible fluid turbulence as an example that is representative, practically important, reasonably simple, and exceedingly well-studied. We use direct numerical simulations of freely decaying turbulence in two spatial dimensions to infer an effective field theory comprising explicit, interpretable evolution equations for both the large (resolved) and small (modeled) scales. The resulting closed system of equations is capable of accurately describing the effect of small scales, including backscatter—the flow of energy from small to large scales, which is particularly pronounced in two dimensions—which is an outstanding challenge that, to our knowledge, no existing alternative successfully tackles.
Introduction—Construction of mathematical models of physical systems—at length, time and, energy scales of interest—has long stood at the foundation of physics. This cornerstone paved the way to improving our understanding of the physical world and making numerous groundbreaking predictions. Often, a fundamental theory is available which provides an exact description of the microscopic details but is practically intractable or fails to generate the needed insight at the scales of interest. In such cases, it is desirable to construct an effective field theory that only describes the relevant degrees of freedom at the scales of interest without trying to model microscopic details. Effective theories are commonly used in particle, nuclear [1], and condensed matter physics [2], but examples such as the Chapman-Enskog theory [3] are found in classical physics as well.
A top-down effective theory can be derived formally from the corresponding fundamental theory only in very specific limits, e.g., the Chapman-Enskog theory requires the gas to be near local thermodynamic equilibrium and the Knudsen number (characterizing the mean free path) to be small. However, one can ask whether a sufficiently general approach exists that, at least in some classes of problems, allows an effective theory to be inferred in a systematic way from the corresponding fundamental theory. A particular class of problems where this capability would be most welcome involves multiscale phenomena, such as turbulence in fluids and plasmas, which can feature structures with length and time scales varying by tens of orders of magnitudes.
Consider, for instance, Newtonian fluids which are well-described by the continuity and Navier-Stokes (momentum) equations
$\nabla\cdot{\bf u}$ $=0,$ (1a)
$\partial_{t}{\bf u}+{\bf u}\cdot\nabla{\bf u}$ $=-\nabla p+\nu\nabla^{2}{\bf u},$ (1b)
where $\nu$ is the kinematic viscosity, down to the scale of microns. The momentum equation features a nonlinearity which both gives rise to the multiscale nature of turbulent flows [4] and severely complicates analysis. Furthermore, this multiscale nature of turbulence makes solution of the fundamental transport equations via direct numerical simulations (DNS) infeasible in practical applications. In particular, the ratio between the largest and smallest length scales in three-dimensional turbulence scales as $M\propto\mathrm{Re}^{3/4}$, where Re is the Reynolds number, while fully resolved DNS requires solving equations for $M^{3}$ degrees of freedom at each time step. For instance, atmospheric jet flows and oceanic currents are described by Re ranging between $10^{10}$ and $10^{12}$, which corresponds to $10^{22}\lesssim M^{3}\lesssim 10^{27}$.
To circumvent this limitation, various models of fluid turbulence have been introduced which involve either temporal or spatial coarse-graining as in, respectively, the Reynolds-averaged Navier-Stokes (RANS) or large eddy simulation (LES). While the coarse-graining procedure itself is formal and general, the equations describing the averaged fields and their fluctuations cannot be closed without making specific assumptions, yielding a phenomenological description. In the case of both RANS and LES, an additional closure term appears in the momentum equation. It represents the effect of the small (modeled) scales on the large (resolved) ones and its functional form is assumed based on various empirical arguments rather than derived.
Figure 1: Representative flow fields used to generate the test data. Shown is the vorticity field.
For instance, in LES, the velocity and pressure fields are decomposed into the large- and small-scale components, $\phi=\bar{\phi}+\phi^{\prime}$, using the filtering operator
$\overline{\phi}\equiv\mathcal{F}(\phi)\equiv\int d\textbf{r}^{\prime}\>G_{\Delta}(\textbf{r}-\textbf{r}^{\prime})\phi(r^{\prime}),$ (2)
where $\Delta$ is the cutoff length scale that defines the resolution of the resulting coarse-grained description. Applying this operator to the fundamental model (1), one obtains a pair of equations for the large-scale components
$\nabla_{i}\bar{u}_{i}$ $=0,$ (3a)
$\partial_{t}\bar{u}_{i}+\bar{u}_{j}\nabla_{j}\bar{u}_{i}$ $=-\nabla_{i}\bar{p}+\nu\nabla^{2}\bar{u}_{i}-\nabla_{j}\tau_{ij},$ (3b)
where we use Einstein implicit summation notation. The closure term $\nabla_{j}\tau_{ij}$ is expressed via the subgrid-scale (SGS) stress tensor $\tau_{ij}=\overline{u_{i}u_{j}}-\bar{u}_{i}\bar{u}_{j}$, which generally cannot be expressed in terms of the resolved fields (here, $\bar{\bf u}$ and $\bar{p}$) and thus requires modeling.
A variety of phenomenological models, each with its strengths and weaknesses, have been proposed. Among the most widely used models are those based on the eddy viscosity hypothesis such as the Smagorinsky model [5] and its dynamic version [6, 7]. Another popular type are the similarity models [8, 9]. Models of both types, as well as hybrids thereof (e.g., the dynamic mixed model [8]) perform well on some benchmarks [10]. However, they are all based on physical assumptions—such as statistical homogeneity, isotropy, and scale invariance—and struggle when these assumptions break, e.g., in the presence of coherent structures. Recent development of data-driven approaches has led to a flurry of efforts to construct closures as functions of the resolved fields by leveraging high-fidelity DNS data. However, only a small fraction of those studies were aimed at obtaining an explicit functional form of the relation, as opposed to black-box relations encoded by neural networks (see, e.g., [11]). Examples include approaches based on sparse regression (e.g., random forest regression [12], sequential thresholding ridge regression [13], relevance vector machine [14]), symbolic regression (e.g., genetic expression programming [15, 16]) or a combination of symbolic and sparse regression [17]. However, none of the data-driven models outperform the leading phenomenological models.
Coherent structures can be very prominent in some types of turbulent flows, e.g, those that are effectively two-dimensional. Such flows often feature strong local fluxes of energy from small to large scales. For a variety of reasons, no existing SGS models—phenomenological or data-driven—are capable of correctly describing backscatter, which plays a key role in a variety of physical phenomena. One prominent example is transport of angular momentum in accretion disks, where magneto-rotational instability that takes place at small scales crucially affects the large scales. Although efforts to develop an LES-like description of scale interaction in the context of magnetohydrodynamics have been made [18], proper treatment of backscatter in this and other contexts remains an open problem.
This letter introduces a general data-driven framework for inferring an effective field theory of multiscale phenomena from the underlying fundamental description. This framework requires no specific phenomenological assumptions and ensures that the resulting theory is equivariant and interpretable. Unlike the bulk of SGS models, the dynamics of small scales and scale interaction are characterized explicitly, which enables an accurate description of backscatter in the case of fluid turbulence.
Methodology—We will retain the coarse-graining approach used in LES to separate the degrees of freedom with length scales larger and smaller than an arbitrary cutoff scale $\Delta$. The choice of the kernel $G$ in equation (2) is partly arbitrary but must satisfy a number of constraints such as locality in physical and Fourier space [19, 20]. We use a normalized Gaussian kernel (with second moment $\sigma^{2}=\Delta^{2}/12$) exclusively in this letter as it maximizes locality in both spaces. To make the description finite-dimensional (e.g., to enable numerical solution), we additionally apply a sharp spectral filter, which cuts off scales past $\Delta$, making the combined filter non-invertible. The resulting splitting formally defines the governing equations for both large scales (here the system (3)) and small scales. The variables describing small scales and the associated governing equations are discarded in LES—which leads to the problems discussed previously—but are occasionally retained in RANS. In the latter case, the effect of small scales is represented by an SGS stress tensor whose formal evolution equation [21] includes variables other than $\bar{u}_{i}$, $\bar{p}$ and $\tau_{ij}$, generated by the nonlinearity, so that the resulting system of equations is not closed.
Construction of a closed system of equations requires identifying (1) proper variables describing small scales, (2) approximate governing equations for these variables, and (3) an approximate constitutive relation for the SGS tensor. While it is natural to expect the small scales to be described by additional tensor field(s) with their own evolution equation(s), the choice of these fields is not obvious a priori. The SGS tensor itself is generally a poor choice, as it also contains substantial contributions from the large scales. This can be seen easily by considering the exact Galilean-invariant decomposition [22, 23]
$\tau_{ij}$ $=$ $L_{ij}+C_{ij}+R_{ij}$ (4a)
$L_{ij}$ $=$ $\overline{\overline{u}_{i}\overline{u}_{j}}-\overline{\overline{u}}_{i}\overline{\overline{u}}_{j}$ (4b)
$C_{ij}$ $=$ $\overline{\overline{u}_{i}u^{\prime}_{j}+u^{\prime}_{i}\overline{u_{j}}}-\overline{\overline{u}}_{i}\overline{u^{\prime}}_{j}-\overline{u^{\prime}}_{i}\overline{\overline{u_{j}}}$ (4c)
$R_{ij}$ $=$ $\overline{u^{\prime}_{i}u^{\prime}_{j}}-\overline{u^{\prime}}_{i}\overline{u^{\prime}}_{j}$ (4d)
of the SGS tensor into the Leonard ($L$), Reynolds ($R$), and cross ($C$) stress tensors describing, respectively, the interaction of large scales, interaction of small scales, and interscale interaction. While both $\bf u$} and $\bf u$$^{\prime}$ (and thereby $L$, $C$ and $R$) can be reconstructed from $\bar{\bf u}$ exactly for the Gaussian filter, the sharp spectral filter is not invertible and so some of the tensors may not be accurately described in terms of the resolved fields.
Both the governing equations for the small-scale variables and the constitutive relation for the SGS stress tensor can be inferred using a data-driven approach. Specifically, we use the SPIDER framework [24] which has been thoroughly validated by recovering fundamental models describing active nematics [25], turbulent channel flow [26], and turbulent convection [27]. SPIDER combines group representation theory, weak formulation of differential equations, and sparse regression to infer complete equivariant models, including boundary conditions, directly from data (experimental or numerical).
SPIDER searches for sparse relations explicitly equivariant with respect to a relevant symmetry group. This is done systematically and comprehensively by exploring search spaces corresponding to different irreducible representations. For instance, for the rotation group, each search space is defined by a basis set of tensors $\{F^{1}_{k},F^{2}_{k},\cdots\}$ that transform according to an irreducible representation $k$ and are constructed using contractions and tensor products of the variables (e.g., $\overline{u}_{i},\overline{p},\cdots$) and their derivatives (e.g., $\partial_{t}\overline{u}_{i},\nabla_{i}\overline{p},\nabla_{j}\overline{u}_{i},\cdots$). The tensor libraries $\mathcal{L}_{k}=\{F^{1}_{k},\cdots,F^{N_{k}}_{k}\}$ are then finitely truncated by selecting an upper bound on the complexity of terms, typically defined as the number of variables and differential operators appearing in a given term. In this way, multiple physical relations can be identified in the form
$c_{1}F^{1}+\cdots+c_{N}F^{N}=0,$ (5)
where $c_{n}$ are learned constant coefficients.
In this letter, we consider non-relativistic flows, so the relevant symmetries include translations, rotations, and Galilean transformations. For simplicity, we consider only the integer-rank tensor (not spinor) representations of the rotation group. Evaluating equation (5) in weak form over a collection of spatiotemporal regions yields an overdetermined linear system $Q{\bf c}=0$ which is solved to find an approximate sparse solution ${\bf c}=(c_{1},\cdots,c_{N}$). Depending on the objective, one can either require $|{\bf c}|=1$ (homogeneous regression) or $c_{m}=1$ for some $1\leq m\leq N$ (inhomogeneous regression). Finally, the scaling of the coefficients with parameters (here, $\Delta$ and $\nu$) is inferred through dimensional analysis and verified by generating data for a range of parameters.
We consider two-dimensional turbulence, which features pronounced coherent structures (eddies) that break statistical isotropy/homogeneity as well as strong backscatter [28], making it an ideal testing ground. We use a pseudospectral solver [11] to generate fully resolved DNS data on a square domain of size $\ell=2\pi$ with periodic boundary conditions. Figure 1 shows snapshots of such flows representing the inverse cascade (F1), freely decaying turbulence (F2) and the direct cascade (F3) used as initial conditions. For each of the three initial conditions, a numerical solution corresponding to $O(1)$ eddy turnover time is generated for each value of the viscosity corresponding to Reynolds number $Re=u_{rms}\ell/\nu$ equal to $10^{4}$, $10^{5}$, and $10^{6}$ on computational grids with respective resolutions of $2048^{2}$, $4096^{2}$, and $4096^{2}$. Next, coarse-grained datasets are generated by applying the filtering operator for a range of cutoff scales $\Delta$. Finally, equations are inferred from the coarse-grained data using the Python implementation of SPIDER [24].
Results—To determine whether the SGS stress tensor can be represented by functions of the resolved fields $\bar{\bf u}$ and $\bar{p}$, we use them to construct the respective rank-0 and symmetric trace-free rank-2 libraries and employ inhomogeneous regression. Across all data sets we consistently find
$\tau_{ij}$ $\approx c_{1}\nabla_{k}\bar{u}_{i}\nabla_{k}\bar{u}_{j}+c_{2}\nabla_{k}\nabla_{m}\bar{u}_{i}\nabla_{k}\nabla_{m}\bar{u}_{j},$ (6)
where the second right-hand-side term is generally small compared to the first one. Incidentally, this result corresponds to the formal series expansions of $L+C$ in powers of $\Delta$ [20]. Using dimensional analysis and the numerical values of the coefficients identified by SPIDER, we find $c_{1}=\Delta^{2}/12$ and $c_{2}=\Delta^{4}/288$, consistent with the coefficients of the formal expansion.
Figure 2: Representative snapshot of the energy flux $\Pi$ for DNS (a), the LCR model (b), DS model (c), and DM model (d) for the flow F2 at $Re\sim 10^{6}$ and cutoff scale $\Delta=\ell/64$.
On the other hand, the tensor $R$, which is generally small compared with $L$ and $C$, cannot be expressed in terms of $\bar{\bf u}$ or $\bar{p}$ with even moderate accuracy. Therefore, we simply treat $R$ as an additional variable in the effective field theory, one that describes small scales, and parameterize the SGS tensor as
$\tau_{ij}\approx\frac{\Delta^{2}}{12}\nabla_{k}\bar{u}_{i}\nabla_{k}\bar{u}_{j}+\frac{\Delta^{4}}{288}\nabla_{k}\nabla_{m}\bar{u}_{i}\nabla_{k}\nabla_{m}\bar{u}_{j}+R_{ij}$ (7)
which we refer to as the LCR model. This model not only reproduces the individual components of the SGS stress tensor faithfully but also accurately captures the spatiotemporal structure of the energy flux $\Pi\equiv-\tau_{ij}\,\bar{S}_{ij}$ (where $S_{ij}$ denotes the symmetric part of $\nabla_{i}u_{j}$) including backscatter (regions where $\Pi<0$). This is illustrated in figure 2 for the freely decaying flow F2. For reference, the figure also shows the predictions of the widely used dynamic Smagorinsky (DS) model as well as the dynamic mixed (DM) model (see End Matter), which is considered to be one of the most accurate LES models [29]. Both phenomenological models fail to predict local fluxes.
The inclusion of $R$ as a new field requires an additional governing equation in order to close the system. This equation can be inferred using rank-0 and symmetric trace-free rank-2 libraries constructed from $\bar{u}_{i}$, $\bar{p}$, and $R_{ij}$. We robustly discover an evolution equation of the form
$\partial_{t}R_{ij}$ $=c_{1}\bar{u}_{k}\nabla_{k}R_{ij}+c_{2}(R_{ik}\nabla_{k}\bar{u}_{j}+R_{jk}\nabla_{k}\bar{u}_{i})$
$+c_{3}\nabla^{2}R_{ij}+c_{4}R_{ij}.$ (8)
Using a combination of dimensional analysis and numerical values of the coefficients for different data sets, we find $c_{1}=-1$, $c_{2}=1$ and $c_{3}=\nu$. For $c_{4}$, we find that it has both dimensions and numerical values consistent with the inverse of the eddy turnover time, exhibiting no clear dependence on the parameters $\nu$ and $\Delta$. Therefore, we relax the assumption of translational invariance made by SPIDER that requires $c_{4}$ to be independent of position and time and take it to be a function of the resolved flow. For instance, in the inviscid limit, rescaling the velocity of the initial condition by a constant factor changes the inverse time scale and, hence, $c_{4}$ by the same factor, so $c_{4}$ should be linear in $\bar{u}$. Because $c_{4}$ is a scalar, it should be expressible as a function of the energy $E=u_{i}u_{i}/2$ as well as the invariants of the tensor $\nabla_{i}\bar{u}_{j}$ [30]; in 2D, these are given by $|\bar{S}|=\sqrt{2\bar{S}_{ij}\bar{S}_{ij}}$ and $|\bar{\Omega}|=\sqrt{2\bar{\Omega}_{ij}\bar{\Omega}_{ij}}$, where ${\Omega}_{ij}$ is the antisymmetric component of $\nabla_{i}u_{j}$. Dimensional arguments would then require $c_{4}=\Delta^{-1}\sqrt{E}f(\Delta|\bar{\Omega}|/\sqrt{E},\Delta|\bar{S}|/\sqrt{E},\cdots)$, where $f$ is an arbitrary function. A general relation of this form can be inferred from the data using, e.g., symbolic regression. However, a good approximation can be found using a series expansion of $f$:
$c_{4}=\alpha_{1}|\bar{S}|+\alpha_{2}|\bar{\Omega}|+\alpha_{3}\Delta^{-1}\sqrt{E}+\cdots,$ (9)
where $\alpha_{j}$ are non-dimensional coefficients. Sparse regression yields $c_{4}\approx-|\bar{S}|/2$. Collecting everything, we find
$\partial_{t}R_{ij}$ $+\bar{u}_{k}\nabla_{k}R_{ij}=R_{ik}\nabla_{k}\bar{u}_{j}+R_{jk}\nabla_{k}\bar{u}_{i}$ (10)
$+\nu\nabla^{2}R_{ij}-\frac{1}{2}|\bar{S}|R_{ij}$
To measure the accuracy in predicting a tensor variable, we will employ a magnitude-aware correlation
$\mathcal{C}(A,B)=\frac{\langle A_{ij\cdots}B_{ij\cdots}\rangle}{\max\left(\langle A_{ij\cdots}A_{ij\cdots}\rangle,\langle B_{ij\cdots}B_{ij\cdots}\rangle\right)}$ (11)
between tensor fields $A$ (ground truth) and $B$ (prediction), where $\langle\cdot\rangle$ denotes spatial averaging, with $\mathcal{C}(A,B)=1$ only when $A=B$. Using $C_{R}=\mathcal{C}(\partial_{t}R_{\text{DNS}},\partial_{t}R_{\text{LES}})$ as a metric, we find equation (10) to accurately describe all flow regimes for sufficiently small $\Delta$ (cf. figure 3). As expected, the accuracy deteriorates as $\Delta$ approaches the scale of coherent structures, since the coarse-grained description is not meant to model these.
Figure 3: The accuracy of the evolution equation (10), quantified by the correlation $C_{R}$, as a function of $\Delta/\ell$ for F1 (blue), F2 (orange) and F3 (yellow) at $Re=10^{6}$.
Conclusions—Taken together, the closed system of equations (3), (7) and (10) defines an effective field theory which describes incompressible fluid turbulence up to the cutoff scale $\Delta$. By construction, this theory is explicitly equivariant with respect to translations and rotations but also turns out to be explicitly Galilean-invariant, as it should. Not only is the resulting description more accurate and general than state-of-the-art LES models, it is explicit and easily interpretable, which allows one to make specific predictions. For example, at scales $\Delta$ small compared to the size of large-scale coherent structures, the energy flux is dominated by the contribution from the $C$ tensor and scales as $\Delta^{4}$, which is not directly implied by the fundamental model (1).
It is worth emphasizing that the data-driven framework described in this letter requires a minimal number of ingredients—the fundamental model (which could also be derived in a data-driven way [26]), some of the fundamental symmetries of the problem, and dimensional analysis—and no specific phenomenological assumptions. Hence, it should readily generalize to other multiscale problems with polynomial nonlinearities.
Acknowledgments—This work was supported by the Defense Advanced Research Projects Agency (DARPA).

### v2

@@
 A top-down effective theory can be derived formally from the corresponding fundamental theory only in very specific limits, e.g., the Chapman-Enskog theory requires the gas to be near local thermodynamic equilibrium and the Knudsen number (characterizing the mean free path) to be small.
-However, one can ask whether a sufficiently general approach exists that, at least in some classes of problems, allows an effective theory to be inferred in a systematic way from the corresponding fundamental theory.
+However, when formal derivations fail, one can ask whether a sufficiently general approach exists that, at least in some classes of problems, allows an effective theory to be inferred in a systematic way from the corresponding fundamental theory.
 A particular class of problems where this capability would be most welcome involves multiscale phenomena, such as turbulence in fluids and plasmas, which can feature structures with length and time scales varying by tens of orders of magnitudes.
@@
 For instance, atmospheric jet flows and oceanic currents are described by Re ranging between $10^{10}$ and $10^{12}$, which corresponds to $10^{22}\lesssim M^{3}\lesssim 10^{27}$.
-To circumvent this limitation, various models of fluid turbulence have been introduced which involve either temporal or spatial coarse-graining as in, respectively, the Reynolds-averaged Navier-Stokes (RANS) or large eddy simulation (LES).
+To circumvent this limitation, various models of fluid turbulence have been introduced which involve either temporal or spatial coarse-graining as in the Reynolds-averaged Navier-Stokes (RANS) or large eddy simulation (LES), respectively.
 While the coarse-graining procedure itself is formal and general, the equations describing the averaged fields and their fluctuations cannot be closed without making specific assumptions, yielding a phenomenological description.
@@
 It represents the effect of the small (modeled) scales on the large (resolved) ones and its functional form is assumed based on various empirical arguments rather than derived.
-Figure 1: Representative flow fields used to generate the test data. Shown is the vorticity field.
 For instance, in LES, the velocity and pressure fields are decomposed into the large- and small-scale components, $\phi=\bar{\phi}+\phi^{\prime}$, using the filtering operator
@@
 However, none of the data-driven models outperform the leading phenomenological models.
+Figure 1: Representative flow fields used to generate the test data. Shown is the initial vorticity field $\omega$ .
 Coherent structures can be very prominent in some types of turbulent flows, e.g, those that are effectively two-dimensional.
@@
 This letter introduces a general data-driven framework for inferring an effective field theory of multiscale phenomena from the underlying fundamental description.
-This framework requires no specific phenomenological assumptions and ensures that the resulting theory is equivariant and interpretable.
-Unlike the bulk of SGS models, the dynamics of small scales and scale interaction are characterized explicitly, which enables an accurate description of backscatter in the case of fluid turbulence.
-Methodology—We will retain the coarse-graining approach used in LES to separate the degrees of freedom with length scales larger and smaller than an arbitrary cutoff scale $\Delta$.
+This framework requires no phenomenological assumptions and ensures that the resulting theory is equivariant and interpretable.
+To illustrate this framework, we apply it to fluid turbulence and infer a subgrid-scale model that accurately describes backscatter.
+This is due, in no small part, to the model’s explicit characterization of the subgrid scales.
+Methodology—We will retain the coarse-graining approach used in LES to separate the degrees of freedom with length scales larger and smaller than a cutoff scale $\Delta$; while $\Delta$ is an adjustable parameter of the model, it must be at a scale substantially smaller than the scale of coherent structures, as SGS models are not designed to capture these.
 The choice of the kernel $G$ in equation (2) is partly arbitrary but must satisfy a number of constraints such as locality in physical and Fourier space [19, 20].
@@
 The resulting splitting formally defines the governing equations for both large scales (here the system (3)) and small scales.
-The variables describing small scales and the associated governing equations are discarded in LES—which leads to the problems discussed previously—but are occasionally retained in RANS.
+The variables describing small scales and their associated governing equations are discarded in LES—which leads to the problems discussed previously—but are occasionally retained in RANS.
 In the latter case, the effect of small scales is represented by an SGS stress tensor whose formal evolution equation [21] includes variables other than $\bar{u}_{i}$, $\bar{p}$ and $\tau_{ij}$, generated by the nonlinearity, so that the resulting system of equations is not closed.
@@
 of the SGS tensor into the Leonard ($L$), Reynolds ($R$), and cross ($C$) stress tensors describing, respectively, the interaction of large scales, interaction of small scales, and interscale interaction.
-While both $\bf u$} and $\bf u$$^{\prime}$ (and thereby $L$, $C$ and $R$) can be reconstructed from $\bar{\bf u}$ exactly for the Gaussian filter, the sharp spectral filter is not invertible and so some of the tensors may not be accurately described in terms of the resolved fields.
-Both the governing equations for the small-scale variables and the constitutive relation for the SGS stress tensor can be inferred using a data-driven approach.
-Specifically, we use the SPIDER framework [24] which has been thoroughly validated by recovering fundamental models describing active nematics [25], turbulent channel flow [26], and turbulent convection [27].
-SPIDER combines group representation theory, weak formulation of differential equations, and sparse regression to infer complete equivariant models, including boundary conditions, directly from data (experimental or numerical).
-SPIDER searches for sparse relations explicitly equivariant with respect to a relevant symmetry group.
-This is done systematically and comprehensively by exploring search spaces corresponding to different irreducible representations.
-For instance, for the rotation group, each search space is defined by a basis set of tensors $\{F^{1}_{k},F^{2}_{k},\cdots\}$ that transform according to an irreducible representation $k$ and are constructed using contractions and tensor products of the variables (e.g., $\overline{u}_{i},\overline{p},\cdots$) and their derivatives (e.g., $\partial_{t}\overline{u}_{i},\nabla_{i}\overline{p},\nabla_{j}\overline{u}_{i},\cdots$).
-The tensor libraries $\mathcal{L}_{k}=\{F^{1}_{k},\cdots,F^{N_{k}}_{k}\}$ are then finitely truncated by selecting an upper bound on the complexity of terms, typically defined as the number of variables and differential operators appearing in a given term.
-In this way, multiple physical relations can be identified in the form
-$c_{1}F^{1}+\cdots+c_{N}F^{N}=0,$ (5)
-where $c_{n}$ are learned constant coefficients.
-In this letter, we consider non-relativistic flows, so the relevant symmetries include translations, rotations, and Galilean transformations.
-For simplicity, we consider only the integer-rank tensor (not spinor) representations of the rotation group.
-Evaluating equation (5) in weak form over a collection of spatiotemporal regions yields an overdetermined linear system $Q{\bf c}=0$ which is solved to find an approximate sparse solution ${\bf c}=(c_{1},\cdots,c_{N}$).
-Depending on the objective, one can either require $|{\bf c}|=1$ (homogeneous regression) or $c_{m}=1$ for some $1\leq m\leq N$ (inhomogeneous regression).
-Finally, the scaling of the coefficients with parameters (here, $\Delta$ and $\nu$) is inferred through dimensional analysis and verified by generating data for a range of parameters.
-We consider two-dimensional turbulence, which features pronounced coherent structures (eddies) that break statistical isotropy/homogeneity as well as strong backscatter [28], making it an ideal testing ground.
-We use a pseudospectral solver [11] to generate fully resolved DNS data on a square domain of size $\ell=2\pi$ with periodic boundary conditions.
+While both $\bf u$} and $\bf u$$^{\prime}$ (and thereby $L$, $C$ and $R$) can be reconstructed from $\bar{\bf u}$ exactly for the Gaussian filter, the sharp spectral filter is not invertible and so some of the tensors may not be accurately reconstructed in terms of the resolved fields.
+On a sufficiently fine grid and for kernels with finite moments, any filtered field can be approximated using the moment expansion [20] as a series in $\Delta/\ell_{c}$, where $\ell_{c}$ is a characteristic length scale such as a vortex size.
+In particular, one finds $L=O((\Delta/\ell_{c})^{2})$, $C=O((\Delta/\ell_{c})^{4})$ and $R=O((\Delta/\ell_{c})^{6})$, which means that $L$ provides the dominant contribution to $\tau$ while $C$ and $R$ yield progressively smaller corrections.
+The leading order term of the moment expansion representation of $\tau$ (and $L$) yields the nonlinear gradient model [32] that was recently rediscovered using a data-driven approach [14].
+While this model yields a reasonably accurate approximation of the SGS tensor, it incorrectly predicts that the energy flux $\Pi\equiv-\tau_{ij}\,\bar{S}_{ij}$ (where $S_{ij}$ denotes the symmetric part of $\nabla_{i}u_{j}$) vanishes identically.
+A correct description of the fluxes requires a far more accurate approximation of the SGS tensor.
+We consider two-dimensional turbulence, which features pronounced coherent structures (eddies) that break statistical isotropy/homogeneity and is characterized by strong backscatter [28], making it an ideal testing ground.
+We use a pseudospectral solver jax-cfd [34] to generate fully resolved DNS data on a square domain of size $\ell=2\pi$ with periodic boundary conditions.
 Figure 1 shows snapshots of such flows representing the inverse cascade (F1), freely decaying turbulence (F2) and the direct cascade (F3) used as initial conditions.
-For each of the three initial conditions, a numerical solution corresponding to $O(1)$ eddy turnover time is generated for each value of the viscosity corresponding to Reynolds number $Re=u_{rms}\ell/\nu$ equal to $10^{4}$, $10^{5}$, and $10^{6}$ on computational grids with respective resolutions of $2048^{2}$, $4096^{2}$, and $4096^{2}$.
-Next, coarse-grained datasets are generated by applying the filtering operator for a range of cutoff scales $\Delta$.
-Finally, equations are inferred from the coarse-grained data using the Python implementation of SPIDER [24].
-Results—To determine whether the SGS stress tensor can be represented by functions of the resolved fields $\bar{\bf u}$ and $\bar{p}$, we use them to construct the respective rank-0 and symmetric trace-free rank-2 libraries and employ inhomogeneous regression.
-Across all data sets we consistently find
-$\tau_{ij}$ $\approx c_{1}\nabla_{k}\bar{u}_{i}\nabla_{k}\bar{u}_{j}+c_{2}\nabla_{k}\nabla_{m}\bar{u}_{i}\nabla_{k}\nabla_{m}\bar{u}_{j},$ (6)
-where the second right-hand-side term is generally small compared to the first one.
-Incidentally, this result corresponds to the formal series expansions of $L+C$ in powers of $\Delta$ [20].
-Using dimensional analysis and the numerical values of the coefficients identified by SPIDER, we find $c_{1}=\Delta^{2}/12$ and $c_{2}=\Delta^{4}/288$, consistent with the coefficients of the formal expansion.
-Figure 2: Representative snapshot of the energy flux $\Pi$ for DNS (a), the LCR model (b), DS model (c), and DM model (d) for the flow F2 at $Re\sim 10^{6}$ and cutoff scale $\Delta=\ell/64$.
-On the other hand, the tensor $R$, which is generally small compared with $L$ and $C$, cannot be expressed in terms of $\bar{\bf u}$ or $\bar{p}$ with even moderate accuracy.
-Therefore, we simply treat $R$ as an additional variable in the effective field theory, one that describes small scales, and parameterize the SGS tensor as
-$\tau_{ij}\approx\frac{\Delta^{2}}{12}\nabla_{k}\bar{u}_{i}\nabla_{k}\bar{u}_{j}+\frac{\Delta^{4}}{288}\nabla_{k}\nabla_{m}\bar{u}_{i}\nabla_{k}\nabla_{m}\bar{u}_{j}+R_{ij}$ (7)
-which we refer to as the LCR model.
-This model not only reproduces the individual components of the SGS stress tensor faithfully but also accurately captures the spatiotemporal structure of the energy flux $\Pi\equiv-\tau_{ij}\,\bar{S}_{ij}$ (where $S_{ij}$ denotes the symmetric part of $\nabla_{i}u_{j}$) including backscatter (regions where $\Pi<0$).
-This is illustrated in figure 2 for the freely decaying flow F2.
-For reference, the figure also shows the predictions of the widely used dynamic Smagorinsky (DS) model as well as the dynamic mixed (DM) model (see End Matter), which is considered to be one of the most accurate LES models [29].
-Both phenomenological models fail to predict local fluxes.
+For each of the three initial conditions, a numerical solution corresponding to about one eddy turnover time is generated on a computational grid with resolution of $2048^{2}$, $4096^{2}$, and $4096^{2}$ for $\nu=10^{-4}$, $10^{-5}$, and $10^{-6}$, respectively.
+The corresponding Reynolds numbers $Re=u_{\rm RMS}\ell/\nu$ are reported in table 1.
+The initial conditions and the software for generating the data are available online [36].
+Next, the numerical solutions are coarse-grained by applying the filtering operator over a range of cutoff scales $\Delta$.
+Afterwards, the functional form of the SGS model is inferred from the coarse-grained data using the Python implementation of SPIDER [24, 35].
+SPIDER combines group representation theory (to construct term libraries which define the search spaces), weak formulation of differential equations (to reduce symbolic relations involving derivatives to a system of linear equations for a set of coefficients), and sparse regression (to infer equivariant functional relations corresponding to different irreducible representations of the relevant symmetry group, in this case, SO(2)) [25, 26, 27].
+Finally, the relationship between the parameters of the effective field theory and the fundamental description (e.g., $\nu$) as well as the filter scale $\Delta$ are determined using a combination of dimensional analysis and scaling arguments.
+Results—To determine whether the SGS stress tensor can be represented by functions of the resolved fields $\bar{\bf u}$ and $\bar{p}$, we use them to construct the respective rank-0 and symmetric trace-free rank-2 term libraries and employ inhomogeneous regression.
+Inhomogeneous regression searches for sparse relations in the form $\tau=f(\bar{\bf u},\bar{p})$ with the known left-hand-side.
+Across all data sets, we consistently find the nonlinear gradient model (NGM2)
+$\tau_{ij}$ $\approx c_{1}(\nabla_{k}\bar{u}_{i})(\nabla_{k}\bar{u}_{j})\equiv\tau_{ij}^{(2)}.$ (5)
+Using dimensional analysis and the numerical values of the coefficients identified by SPIDER for different data sets, we find $c_{1}=\Delta^{2}/12$, consistent with the moment expansion of $L$.
+To obtain a more accurate approximation, we repeat the inhomogeneous regression for the difference $\tau_{ij}-\tau_{ij}^{(2)}$, which consistently yields the next order correction
+$\tau_{ij}-\tau_{ij}^{(2)}$ $\approx c_{2}(\nabla_{k}\nabla_{m}\bar{u}_{i})(\nabla_{k}\nabla_{m}\bar{u}_{j})\equiv\tau_{ij}^{(4)}.$ (6)
+Dimensional analysis and scaling yield $c_{2}=\Delta^{4}/288$, consistent with the moment expansion of $L+C-\tau^{(2)}$.
+Note that, for small $\Delta$, $\tau^{(4)}$ becomes small compared with $\tau^{(2)}$ and, as a result, is generally omitted by inhomogeneous regression for $\tau$.
+The iterative procedure described here reliably picks up this term.
+We will refer to the parameterization $\tau=\tau^{(2)}+\tau^{(4)}$ as the NGM4 model.
+Figure 2: The energy flux $\Pi$ describing the flow F2 shown in Figure 1(b) at $\nu=10^{-6}$ and cutoff scale $\Delta=\ell/64$ for DNS (a), the NGMR model (b), DS model (c), and DM model (d).
+All fields were calculated from DNS data.
+As we show later, NGM4 is still not capable of accurately describing the energy flux.
+Hence, we repeat the iterative procedure to identify the next, even smaller correction $\tau^{(6)}=O(\Delta^{6})$.
+However, this time we find no robust parameterization for $\tau^{(6)}$ in terms of $\bar{\bf u}$ and $\bar{p}$ with even moderate accuracy, which means that $\tau^{(6)}$ must be expressed in terms of an additional tensor field describing the small scales.
+Since NGM4 already accounts for the contributions from $L$ and $C$ in the decomposition (4a), it is natural to assume that $\tau^{(6)}\approx R$ and treat $R$ as an additional variable in the effective field theory.
+This leads to the following parameterization of the SGS tensor
+$\tau\approx\tau^{(2)}+\tau^{(4)}+R$ (7)
+which we refer to as the NGMR model.
+This model not only reproduces the individual components of the SGS stress tensor faithfully but also accurately captures the spatiotemporal structure of the energy flux $\Pi$ including backscatter (regions where $\Pi<0$) as illustrated in figure 2 for a particular flow and a particular choice of viscosity and the filter scale.
+For reference, the figure also shows the predictions of the widely used dynamic Smagorinsky (DS) model as well as the dynamic mixed (DM) model (see SI), which is considered to be one of the most accurate LES models [29].
+In fact, the NGMR model succeeds on every metric and in every case we considered, while the phenomenological models rarely do.
+To measure the accuracy of an LES model in predicting a tensor variable, we employ a magnitude-aware correlation
+$\mathcal{C}(A,B)=\frac{\langle A_{ij\cdots}B_{ij\cdots}\rangle}{\max\left(\langle A_{ij\cdots}A_{ij\cdots}\rangle,\langle B_{ij\cdots}B_{ij\cdots}\rangle\right)}$ (8)
+between tensor fields $A$ (ground truth) and $B$ (prediction), where $\langle\cdot\rangle$ denotes spatial averaging, with $\mathcal{C}(A,B)=1$ precisely when $A=B$.
+In particular, $C_{\tau}=C(\tau_{\rm DNS},\tau_{\rm LES})$ and $C_{\Pi}=C(\Pi_{\rm DNS},\Pi_{\rm LES})$ describe the accuracy in predicting the components of the SGS tensor and the energy flux, respectively.
+The accuracy in predicting the net energy flux can be described in terms of the ratio $q_{\Pi}=\langle\Pi_{\text{LES}}\rangle/\langle\Pi_{\text{DNS}}\rangle$.
+All three quantities should be equal to unity (100%) if the prediction is perfect.
+As table 1 shows, both phenomenological models fail to predict the local fluxes for any flows and the net flux for flows F1 and F3.
+An unexpected observation is that a very accurate prediction of $\tau$ (as in the case of NGM4) does not guarantee accuracy in predicting the local and global energy fluxes.
+NGMR is the only model capable of predicting all these quantities across the board.
+$\nu$ $10^{-4}$ $10^{-5}$ $10^{-6}$
+ReRe $2\times 10^{4}$ $2\times 10^{4}$ $3\times 10^{5}$ $3\times 10^{5}$ $2\times 10^{5}$ $3\times 10^{6}$ $2\times 10^{6}$ $2\times 10^{6}$ $3\times 10^{7}$
+model F1 F2 F3 F1 F2 F3 F1 F2 F3
+DS 0.13% 0.01% 0.01% 0.24% 0.01% 0.01% 0.27% 0.01% 0.01%
+DM 82.28% 99.71% 99.67% 73.52% 96.25% 99.67% 71.44% 96.17% 99.66%
+$C_{\tau}$ NGM2 92.27% 99.27% 99.94% 89.21% 99.15% 99.93% 88.44% 99.14% 99.934%
+NGM4 97.70% 99.93% 99.99% 95.94% 99.89% 99.99% 95.40% 99.89% 99.99%
+NGMR 99.47% 99.98% 99.99% 99.34% 99.97% 99.99% 99.33% 99.97% 99.99%
+DS 26.01% 24.86% 22.77% 26.28% 27.87% 24.05% 22.48% 30.63% 24.06%
+DM 27.62% 18.00% 15.64% 29.63% 20.20% 13.41% 27.03% 21.08% 13.40%
+$C_{\Pi}$ NGM2 0% 0% 0% 0% 0% 0% 0% 0% 0%
+NGM4 83.49% 90.57% 99.44% 75.41% 83.81% 94.12% 72.26% 83.07% 93.98%
+NGMR 96.85% 98.02% 99.85% 96.97% 97.52% 98.89% 97.12% 97.40% 98.83%
+DS 131.76% 107.25% 31.53% 129.05% 106.27% 26.87% 128.74% 106.17% 26.86%
+DM 134.12% 109.92% 35.03% 131.36% 108.93% 30.19% 131.04% 108.83% 30.18%
+$q_{\Pi}$ NGM2 0% 0% 0% 0% 0% 0% 0% 0% 0%
+NGM4 59.31% 69.83% 82.33% 49.13% 62.29% 58.60% 46.33% 60.78% 56.28%
+NGMR 104.18% 104.63% 103.38% 103.81% 104.08% 104.94% 102.71% 103.93% 104.63%
+Table 1: The accuracy, averaged in time over the dataset, of the SGS stresses and corresponding energy fluxes for flows F1–F3, with $R$ computed from DNS.
+To test the limits of the inferred parameterization, we purposely chose a coarse resolution with a cutoff scale $\Delta=\ell/64$ that is close to the size of the eddies in the flow F1.
 The inclusion of $R$ as a new field requires an additional governing equation in order to close the system.
-This equation can be inferred using rank-0 and symmetric trace-free rank-2 libraries constructed from $\bar{u}_{i}$, $\bar{p}$, and $R_{ij}$.
+This equation can be inferred using rank-0 and symmetric trace-free rank-2 libraries constructed from $\bar{u}_{i}$, $\bar{p}$, and $R_{ij}$ and homogeneous regression, which searches for sparse relations in the form $f(R,\bar{\bf u},\bar{p})=0$.
 We robustly discover an evolution equation of the form
-$\partial_{t}R_{ij}$ $=c_{1}\bar{u}_{k}\nabla_{k}R_{ij}+c_{2}(R_{ik}\nabla_{k}\bar{u}_{j}+R_{jk}\nabla_{k}\bar{u}_{i})$
+$\partial_{t}R_{ij}$ $\approx c_{1}\bar{u}_{k}\nabla_{k}R_{ij}+c_{2}(R_{ik}\nabla_{k}\bar{u}_{j}+R_{jk}\nabla_{k}\bar{u}_{i})$
 $+c_{3}\nabla^{2}R_{ij}+c_{4}R_{ij}.$ (9)
 Using a combination of dimensional analysis and numerical values of the coefficients for different data sets, we find $c_{1}=-1$, $c_{2}=1$ and $c_{3}=\nu$.
-For $c_{4}$, we find that it has both dimensions and numerical values consistent with the inverse of the eddy turnover time, exhibiting no clear dependence on the parameters $\nu$ and $\Delta$.
-Therefore, we relax the assumption of translational invariance made by SPIDER that requires $c_{4}$ to be independent of position and time and take it to be a function of the resolved flow.
+For $c_{4}$, we find that it has both dimensions and numerical values consistent with the inverse of the eddy turnover time but exhibits no clear dependence on the parameters $\nu$ and $\Delta$.
+Therefore, we relax SPIDER’s assumption of translational invariance that requires $c_{4}$ to be independent of position and time and take it to be a function of the resolved flow.
 For instance, in the inviscid limit, rescaling the velocity of the initial condition by a constant factor changes the inverse time scale and, hence, $c_{4}$ by the same factor, so $c_{4}$ should be linear in $\bar{u}$.
@@
 where $\alpha_{j}$ are non-dimensional coefficients. Sparse regression yields $c_{4}\approx-|\bar{S}|/2$. Collecting everything, we find
-$\partial_{t}R_{ij}$ $+\bar{u}_{k}\nabla_{k}R_{ij}=R_{ik}\nabla_{k}\bar{u}_{j}+R_{jk}\nabla_{k}\bar{u}_{i}$ (10)
+$\partial_{t}R_{ij}$ $+\bar{u}_{k}\nabla_{k}R_{ij}\approx R_{ik}\nabla_{k}\bar{u}_{j}+R_{jk}\nabla_{k}\bar{u}_{i}$ (11)
 $+\nu\nabla^{2}R_{ij}-\frac{1}{2}|\bar{S}|R_{ij}$
-To measure the accuracy in predicting a tensor variable, we will employ a magnitude-aware correlation
-$\mathcal{C}(A,B)=\frac{\langle A_{ij\cdots}B_{ij\cdots}\rangle}{\max\left(\langle A_{ij\cdots}A_{ij\cdots}\rangle,\langle B_{ij\cdots}B_{ij\cdots}\rangle\right)}$ (11)
-between tensor fields $A$ (ground truth) and $B$ (prediction), where $\langle\cdot\rangle$ denotes spatial averaging, with $\mathcal{C}(A,B)=1$ only when $A=B$.
-Using $C_{R}=\mathcal{C}(\partial_{t}R_{\text{DNS}},\partial_{t}R_{\text{LES}})$ as a metric, we find equation (10) to accurately describe all flow regimes for sufficiently small $\Delta$ (cf.
+Using $C_{R}=\mathcal{C}(\partial_{t}R_{\text{DNS}},\partial_{t}R_{\text{NGMR}})$ as a metric, we find equation (11) to accurately describe all flow regimes for sufficiently small $\Delta$ (cf.
 figure 3).
 As expected, the accuracy deteriorates as $\Delta$ approaches the scale of coherent structures, since the coarse-grained description is not meant to model these.
-Figure 3: The accuracy of the evolution equation (10), quantified by the correlation $C_{R}$, as a function of $\Delta/\ell$ for F1 (blue), F2 (orange) and F3 (yellow) at $Re=10^{6}$.
+Figure 3: The accuracy of the evolution equation (11), quantified by the correlation $C_{R}$, as a function of $\Delta/\ell$ for F1 (blue), F2 (orange) and F3 (yellow) at $\nu=10^{-6}$.
 Conclusions—Taken together, the closed system of equations (3), (7) and (11) defines an effective field theory which describes incompressible fluid turbulence up to the cutoff scale $\Delta$.
 By construction, this theory is explicitly equivariant with respect to translations and rotations but also turns out to be explicitly Galilean-invariant, as it should.
-Not only is the resulting description more accurate and general than state-of-the-art LES models, it is explicit and easily interpretable, which allows one to make specific predictions.
+Not only is the resulting description more accurate and general than state-of-the-art LES models—at least in two spatial dimensions—it is explicit and easily interpretable, which allows one to make specific predictions.
 For example, at scales $\Delta$ small compared to the size of large-scale coherent structures, the energy flux is dominated by the contribution from the $C$ tensor and scales as $\Delta^{4}$, which is not directly implied by the fundamental model (1).
-It is worth emphasizing that the data-driven framework described in this letter requires a minimal number of ingredients—the fundamental model (which could also be derived in a data-driven way [26]), some of the fundamental symmetries of the problem, and dimensional analysis—and no specific phenomenological assumptions.
-Hence, it should readily generalize to other multiscale problems with polynomial nonlinearities.
+Our data-driven framework also identifies two methodological problems with traditional modeling approaches.
+Accurate description of inter-scale fluxes requires introduction of additional fields describing subgrid scales which are missing in all LES models.
+Moreover, these fields must have a tensor structure, unlike the most common RANS models such as $k-\epsilon$, $k-\omega$, and Spalart-Allmaras, which all employ scalar fields.
+The traditional modeling approach has required decades of slow iterative progress and numerous—and often poorly justified—phenomenological assumptions to develop the current generation of LES models.
+In contrast, the framework presented here is both far more efficient and produces more accurate SGS models.
+Most importantly, it requires a minimal number of ingredients: the fundamental model (which could also be derived in a data-driven way [26]), some of the fundamental symmetries of the problem, dimensional analysis, and no specific phenomenological assumptions.
+Hence, it should readily generalize to other multiscale problems described by equations with quadratic nonlinearities, as discussed in the SI.
 Acknowledgments—This work was supported by the Defense Advanced Research Projects Agency (DARPA).
+Supplementary Information
+The accuracy and resolution of DNS data
+Unlike conventional approaches, such as numerical convergence studies, which only quantify the relative accuracy of the numerical solutions obtained at different resolutions, SPIDER provides an absolute measure of the accuracy.
+For each inferred equation, the magnitude of the residual is computed that quantifies how large the error of the equation is (in weak form) compared with the magnitude of individual terms [24].
+In particular, the accuracy of the unfiltered equation (1b) is found to be very close to that of the corresponding filtered equation (3b) for any $\Delta$.
+It is the accuracy of the former that determines whether the data is sufficiently well-resolved to infer an accurate constitutive relation for the SGS stress tensor.
+As table 2 illustrates, for the pseudospectral solver [34], the accuracy quickly improves with increasing resolution.
+For each of the three flows, we chose the resolution at which the accuracy plateaus, indicating that it is limited by factors other than the grid resolution.
+A key limiting factor is the precision of numerical quadratures used in the evaluation of the weak form of each term.
+The quadrature uses a fixed number of grid points and does not scale with the resolution of the data.
+Numerical solutions obtained on a $4096^{2}$ grid at the lowest value of the viscosity considered here have an accuracy (i.e., $O(10^{-10})$ for every flow) that is more than sufficient for the purposes of robustly inferring the same SGS model for different choices of $\Delta$.
+NN F1 F2 F3
+128 $(1\times 10^{-4})$ $(3\times 10^{-3})$ $(1\times 10^{-4})$
+256 $(6\times 10^{-5})$ $(1\times 10^{-3})$ $(2\times 10^{-5})$
+512 $3\times 10^{-6}$ $2\times 10^{-7}$ $3\times 10^{-8}$
+1024 $2\times 10^{-7}$ $1\times 10^{-7}$ $4\times 10^{-10}$
+2048 $3\times 10^{-9}$ $2\times 10^{-9}$ $5\times 10^{-11}$
+4096 $3\times 10^{-10}$ $7\times 10^{-13}$ $7\times 10^{-13}$
+8192 $1\times 10^{-11}$ $7\times 10^{-13}$ $4\times 10^{-11}$
+Table 2: The accuracy of the momentum equation (1b) inferred by SPIDER from pseudospectral numerical data generated on an $N\times N$ grid for the flows F1, F2, and F3 at $\nu=10^{-6}$.
+Parentheses indicate instances where the Euler, rather than Navier-Stokes, equation was inferred, with the viscous term missing.
+Comparison to other SGS models
+To illustrate how our approach and the effective field theory introduced here fit in the mature field of turbulence modeling, we draw some parallels with phenomenological modeling approaches and models and provide a more comprehensive comparison of our results with those of the leading LES models.
+Specifically, we use the dynamic Smagorinsky (DS) model and the dynamic mixed (DM) model, neither of which describes subgrid scales explicitly and assumes the SGS to be a function of large (resolved) scales only.
+The DS model uses the eddy viscosity assumption
+$\tau_{ij}=\nu_{e}\bar{S}_{ij},\quad\nu_{e}=-2\,(c_{\text{smg}}\Delta)^{2}|\bar{S}|$ (12)
+where $c_{\text{smg}}$ is a nondimensional scalar field determined by a dynamic procedure [6]. The DM model is a hybrid between a similarity model and the DS model:
+$\tau_{ij}=c_{\text{\text{sim}}}\left(\overline{\tilde{u}_{i}\tilde{u}_{j}}-\overline{\tilde{u}}_{i}\,\overline{\tilde{u}}_{j}\right)-2\,(c_{\text{smg}}\Delta)^{2}|\tilde{S}|\,\tilde{S}_{ij},$ (13)
+where the tilde and overbar describe filtering at cut-off scales $\Delta$ and $\gamma\Delta$ respectively.
+In this paper, the DM model was employed with $c_{\text{sim}}=\gamma=1$ [29], in which case the first term on the right-hand-side of equation (13) corresponds to the Leonard stress tensor ($L$) and the second to the DS model.
+The DS model, as well as its implementation within the DM model, employs local averaging with a Gaussian kernel at the test filter width, $\gamma\Delta$ [31].
+On the one hand, unlike traditional LES models, NGM2 and NGM4 are both derivable from the fundamental description using a formal moment expansion of the SGS stress tensor $\tau$ and do not rely on any phenomenological assumption.
+On the other hand, like traditional models, both NGM2 and NGM4 use a parameterization entirely in terms of the resolved velocity field $\bar{\bf u}$.
+NGMR is neither derivable from the fundamental description nor can it be parameterized entirely in terms of the resolved fields.
+In this sense its structure is more similar to RANS models which introduce one or more additional fields to describe sub-grid-scale degrees of freedom.
+The fundamental difference is that RANS models use phenomenological assumptions to obtain both the evolution equations for the SGS variables and the parameterization of the SGS stress tensor, while NGMR requires no such assumptions.
+The governing equation (11) for the Reynolds stress tensor, which appears in the NGMR model, has no direct analogues in either LES or RANS literature.
+It does, however, contain several terms found in the evolution equations which appear in RANS models, e.g., SSG [33] and its variants, notably the advection term $\bar{u}_{k}\nabla_{k}R_{ij}$ and the production term(s) $R_{ik}\nabla_{k}\bar{u}_{j}+R_{jk}\nabla_{k}\bar{u}_{i}$.
+In fact, equation (11) can be written in a much simpler form
+$\accentset{\nabla}{R}_{ij}=\nu\nabla^{2}R_{ij}-\frac{1}{2}|\bar{S}|R_{ij},$ (14)
+where
+$\accentset{\nabla}{R}_{ij}=\partial_{t}R_{ij}+\bar{u}_{k}\nabla_{k}R_{ij}-R_{ik}\nabla_{k}\bar{u}_{j}-R_{jk}\nabla_{k}\bar{u}_{i}$ (15)
+is known as the upper-convected or Oldroyd time derivative.
+As table 1 illustrates, the DS model fails to predict SGS stresses with any meaningful accuracy for any flow considered here.
+The DM model is quite accurate for the flow F3 which features the largest eddies, but its accuracy quickly deteriorates as the size of the eddies decreases (flow F2 and especially flow F1).
+The accuracy of the NGM2 model exceeds that of the DM model slightly for every flow considered (similar performance is not unexpected since $\tau\approx L$ for both models), with the NGM4 model bringing further improvement in accuracy.
+The NGMR model is not only the most accurate, but it predicts the SGS stresses essentially perfectly for all of the flows and values of $Re$ considered here, with the correlation exceeding 99% in every case.
+Similarly, NGMR is the only model that consistently captures the spatiotemporal structure of the energy fluxes, with the correlation exceeding 97% in every case (cf.
+table 1), consistent with the results shown in figure 2.
+The NGM2 model is well-known to yield $\Pi=0$, while the accuracy of the NGM4 model trails the NGMR model despite reproducing the SGS stress tensor with high precision.
+This illustrates that high values of $C_{\tau}$ are a necessary but not sufficient condition for the accuracy of a SGS parameterization.
+Both the DS and the DM model completely fail to reproduce the spatiotemporal structure of the energy flux for any of the flows considered here.
+Phenomenological models are commonly designed and tuned to reproduce the net energy dissipation.
+Hence, it is perhaps not surprising that both the DS and the DM model manage to predict the net energy flux for the freely decaying flow (F2) with reasonable accuracy despite completely failing at describing its spatiotemporal structure.
+However, both phenomenological models overpredict the net flux for the inverse cascade (F1) and severely underpredict it for the direct cascade (F3), raising the possibility that the good agreement for the flow F2 is purely coincidental.
+The NGM2 model cannot predict the net flux, as expected, while the NGM4 model consistently captures only 50% to 80% of the net flux.
+Once again, the NGMR model is the only one able to consistently predict the net flux, which illustrates the importance of including the contribution from the Reynolds stress tensor in the parameterization of the SGS stress tensor.
+Generalizations to other systems
+In order to see how our modeling framework can be generalized to other systems, consider a system described by a finite number of tensor fields (tensorial indices suppressed) $q^{(\alpha)}$, ${\alpha}=1,2,\cdots$ with evolution equations
+$\partial_{t}q^{{(\alpha)}}+\sum_{k}N_{k}^{(\alpha)}(q)=S^{(\alpha)}(q),$ (16)
+where $q\equiv\{q^{(1)},q^{(2)},\cdots\}$, $S^{(\alpha)}$ are terms linear in $q$, and $N_{k}^{(\alpha)}$ are quadratic nonlinearities.
+We will assume that $N_{k}^{(\alpha)}=D_{k}^{(\alpha)}(A^{(\alpha)}_{k}B_{k}^{(\alpha)})$, where $D_{k}^{(\alpha)}$ are some differential operators acting on tensor products of linear functions $A_{k}^{(\alpha)}(q^{(\gamma)})$ and $B_{k}^{(\alpha)}(q^{(\beta)})$ with arbitrary $\gamma$ and $\beta$.
+Examples of such nonlinearities are the terms $\nabla\cdot({\bf u}\otimes{\bf u})$ in the momentum equation or $\nabla\times({\bf u}\times{\bf B})$ in the induction equation in magnetohydrodynamics.
+Filtering equations (16) yields
+$\partial_{t}\bar{q}^{(\alpha)}+\sum_{k}N_{k}^{(\alpha)}(\bar{q})=S^{(\alpha)}(\bar{q})-\sum_{k}D_{k}^{(\alpha)}Q_{k}^{(\alpha)}(q,\bar{q}),$
+where $D_{k}^{(\alpha)}Q_{k}^{(\alpha)}$ are the closure terms with closure variables
+$Q_{k}^{(\alpha)}=\overline{A_{k}^{(\alpha)}B_{k}^{(\alpha)}}-\bar{A}_{k}^{(\alpha)}\bar{B}_{k}^{(\alpha)}.$
+Each tensor $Q^{(\alpha)}_{k}$ can be decomposed, similar to Equation (4), into components
+$Q_{k}^{(\alpha)}$ $=$ $L_{k}^{(\alpha)}+C_{k}^{(\alpha)}+R_{k}^{(\alpha)}$
+$L_{k}^{(\alpha)}$ $=$ $\overline{\overline{A}_{k}^{(\alpha)}\overline{B}_{k}^{(\alpha)}}-\overline{\overline{A}}_{k}^{(\alpha)}\overline{\overline{B}}_{k}^{(\alpha)}$
+$C_{k}^{(\alpha)}$ $=$ $\overline{\overline{A}_{k}^{(\alpha)}{B^{\prime}}^{(\alpha)}_{k}+\overline{B}_{k}^{(\alpha)}{A^{\prime}}^{(\alpha)}_{k}}-\overline{\overline{A}}_{k}^{(\alpha)}\overline{B^{\prime}}_{k}^{(\alpha)}-\overline{\overline{B}}_{k}^{(\alpha)}\overline{A^{\prime}}_{k}^{(\alpha)}$
+$R_{k}^{(\alpha)}$ $=$ $\overline{{A^{\prime}}_{k}^{(\alpha)}{B^{\prime}}_{k}^{(\alpha)}}-\overline{A^{\prime}}_{k}^{(\alpha)}\overline{B^{\prime}}_{k}^{(\alpha)}$
+describing, respectively, the interactions between resolved scales, resolved and subgrid scales, and subgrid scales. The gauge symmetry
+$Q_{k}^{(\alpha)}(A_{k}^{(\alpha)}+a,B_{k}^{(\alpha)}+b)=Q_{k}^{(\alpha)}(A_{k}^{(\alpha)},B_{k}^{(\alpha)})$
+where $a$ and $b$ are any spatially uniform tensor fields, is inherited individually by each of the three components.
+In the case of the momentum equation, this symmetry corresponds to Galilean invariance.
+Generally, each of the $L_{k}^{(\alpha)},C_{k}^{(\alpha)},R_{k}^{(\alpha)}$ tensors must be approximated well for $Q_{k}^{(\alpha)}$ to be accurately modeled.
+To this end, we distinguish tensors that are resolvable—those that can be parameterized effectively in terms of the respective $\bar{q}^{(\gamma)}$ and $\bar{q}^{(\beta)}$ with the moment expansion on the coarse grid—and unresolvable—those that cannot.
+This notion of resolvability generalizes to arbitrary systems (with quadratic nonlinearities), as the moment expansion is a formal series expansion and does not rely on the system.
+Moreover the hierarchical scaling $L_{k}^{(\alpha)}=O((\Delta/\ell_{c})^{2}),\ C_{k}^{(\alpha)}=O((\Delta/\ell_{c})^{4}),\ R_{k}^{(\alpha)}=O((\Delta/\ell_{c})^{6})$ holds generally, too.
+$L^{(\alpha)}_{k}$ is a function of resolved scales by definition and hence requires no explicit modeling—one can use either its moment expansion or its definition.
+This is the path conventional SGS models tend to follow while ignoring contributions from $C_{k}^{(\alpha)}$ and $R_{k}^{(\alpha)}$.
+$C_{k}^{(\alpha)}$ and $R_{k}^{(\alpha)}$ generally require modeling, since both components formally depend on the subgrid scales.
+However, it may be the case that either (or both) can be parameterized effectively with the moment expansion.
+The resolvability of each $C_{k}^{(\alpha)}$ and $R_{k}^{(\alpha)}$ can be determined by performing iterative inhomogeneous regression and inspecting the structure of the inferred parameterizations and the associated residuals.
+It is possible for all $L^{(\alpha)}_{k},C^{(\alpha)}_{k},R^{(\alpha)}_{k}$ tensors to be parameterized by the moment expansion with adequate accuracy if the spectra of the respective tensors decay quickly enough.
+In such a case, no additional equations are needed to close the system.
+Note that, even if $R_{k}^{(\alpha)}$ is found to be unresolvable in terms of $\bar{q}^{(\gamma)}$ and $\bar{q}^{(\beta)}$, it may still be possible to express $C_{k}^{(\alpha)}$ in terms of $\bar{q}^{(\gamma)}$, $\bar{q}^{(\beta)}$ and $R_{k}^{(\alpha)}$.
+We should emphasize a practical issue that is critical to the success of data-driven modeling but can be easily overlooked.
+It is of utmost importance that all the variables are projected onto the coarse grid before any functional relations are inferred using SPIDER; it is insufficient to only filter the variables.
+Indeed, the filtering operation is invertible in the spatially continuous case (or on a grid sufficiently fine to resolve the solutions of the fundamental equations).
+Specifically, the moment expansion can be used to recover the original variables—and hence the subgrid scales—from their filtered versions with high accuracy on a fine grid.
+As a consequence of this (as well as the fast scaling of the $C$ and especially $R$ component of closure variables with $\Delta$), testing SGS models on fine grids is rather meaningless and should be conducted on grids substantially coarser than those needed to fully resolve the solutions of the fundamental equations.
+Once the unresolvable tensors are ascertained, these tensors should be treated as new variables in the effective field theory (EFT).
+The corresponding evolution or constitutive equations necessary to close the system can be inferred using homogeneous regression.
+These equations are inferred with specific numerical values for each of the coefficients (i.e., parameters of the EFT).
+A combination of dimensional analysis and scaling can be used to identify the functional relation between the EFT parameters and the parameters of the fundamental equations (16).
+As the example of fluid turbulence illustrates, the coefficients of the inferred evolution/constitutive equations can also depend on various tensor invariants.
+For example, some parameters of the evolution/constitutive equation for $R^{(\alpha)}_{k}$ can depend on the invariants $I^{(\alpha)}_{k,\eta}$ of the tensors $\bar{q}^{(\eta)}$, $\Delta\nabla_{i}\bar{q}^{(\eta)}$, $\Delta^{2}\nabla_{i}\nabla_{j}\bar{q}^{(\eta)}$, etc.
+with $\eta=\gamma,\beta$ corresponding to the particular $k$ and $\alpha$.
+The scaling of the invariants with $\Delta$ shows that only a finite number of these (corresponding to the lowest powers of $\Delta$) need to be considered.
+The functional dependence on these invariants can be inferred using either sparse regression or, failing that, symbolic regression.